Goode Company has the following production data for selected months.

Ending
Work in Process


Month

Beginning
Work in Process

Units
Transferred Out


Units

% Complete as to
Conversion Cost

January 0 36,800 10,700 42 %
March 0 42,700 8,500 70
July 0 46,300 21,700 21


Compute equivalent units of production for materials and conversion costs, assuming materials are entered at the beginning of the process.

Materials

Conversion Costs

January

Entry field with correct answer

Entry field with incorrect answer

March

Entry field with correct answer

Entry field with incorrect answer

July

Entry field with correct answer

Entry field with incorrect answer
Answer #1
Month Materials Conversion Costs
January                              47,500.00                                            41,294.00
March                              51,200.00                                            48,650.00
July                              68,000.00                                            50,857.00
Explanation
Month Materials Conversion Costs
January (36,800 + 10,700) = 47,500 [36,800 + (10,700 x 42%)] = 41,294
March (42,700 + 8,500) = 51,200 [42,700 + (8,500 x 70%)] = 48,650
July (46,300 + 21,700) = 68,000 [46,300 + (21,700 x 21%)] = 50,857